# Protecting Copyrighted Material with Unique Identifiers in Large Language Model Training

### v2

@@ 3.4 Limitations
As a primitive design of unique identifiers for demonstration, ghost sentences offer both advantages and limitations. They are transparent, user-friendly, and statistically trustworthy. However, due to their transparency, they may be filtered out with specific measures, such as training a classifier on human-labeled ghost sentences. This approach, though, is costly and may result in many false positives due to diverse custom prefixes as shown in Figure 1. Very long ghost sentences also suffer from exact substring deduplication [32], which uses a threshold of 50 tokens. Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer [54]. Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found [28] (or Figure 5). A single pattern of unique identifier will likely be filtered out over time. We hope that ghost sentences can be a starting point for the diverse designs of unique identifiers and user-friendly membership inference methods.
@@ 4.1 Experimental Detail
In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning [63, 59] stage. At both the two stages, LLMs can use user data for tuning [5, 58, 60, 61, 11, 34].
Models   For instruction tuning, we adopt the LLaMA serires [60], including OpenLLaMA-3B [19], LLaMA-7B, and LLaMA-13B. For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps. The context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works [59, 34]. We maintain the pre-training batch size the same as TinyLlama-1.1B — 1024 examples. A large batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
@@
Dataset   Webis-TLDR-17 [62] contains 3.7M examples with word lengths under 4096. Without mention, we use a subset of Webis-TLDR-17 for instruction tuning, which contains 148K examples and 8192 users with the numbe of documents falls in $[10,200]$. We term this subset as Webis-148K for convenient. For instruction tuning on Webis-148K, LLMs are required to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. For continuing pre-training, we also utilize the LaMini-Instruction [65] and OpenOrca [38, 45, 35] datasets, which contain 2.6M and 3.5M examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9.8M. All data are shuffled during training.
Evaluation and Metrics   For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test. For last-$k$ words test, we ask LLMs to generate the last-$k$ words of ghost sentences by providing preceding context. A beam search with width 5 is used for generation. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$.
@@
4.2 Perplexity Test
To figure out the average repetition $\mu$ of ghost sentences, we randomly generate 30 different ghost sentences with a word length 10. Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K. During test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them. The recall corresponds to a significant level 0.05 and we choose a critical value like Figure 2. When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance. Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences. With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences. For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100. The perplexity of normal context are roughly the same after fine-tuning.
@@ 4.3 Last-$k$ Words Test
The last-$k$ words test is more straightforward for general users compared to the perplexity test. It only requires chatting with LLMs. However, the last-$k$ words test requires more repetition times of ghost sentences. In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences. For all used training examples, we randomly select $m$ users for the insertion of ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. A few key observations:
  • When $\mu\geq 10$, ghost sentences with a word length around 10 are very likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k. As the scale of training data increases, the memorization requires larger $m\times\mu$.
  • The success rate of memorization is jointly determined by $m$ and $\mu$. Notably, $\mu$ is more critical than $m$. Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (Appendix E).
  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (Appendix F).
  • Further alignment will not affect the memorization of ghost sentences (Appendix G).
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works [8]. Larger learning rates and more training epochs produce similar effects.
@@ 4.3.1 Number and Repetition Times
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2(a), the identification accuracy drops with the increasing data scale. For $16$ sentences with $24$ average repetition time in 1.8M training examples, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last-$1$ word prediction. In this case, documents with ghost sentences only account for 0.02% of all 1777K examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is quite feasible. Especially for content platforms, which can easily achieve such a goal.
@@ 4.3.2 Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 2(b), we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length around 10 words. For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) [22]. It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication [32], which use a threshold of 50 tokens.
@@ 4.3.3 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance. Table 3(a) displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models [8]. This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model [6].
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 3(a). This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a LLaMA-3B model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA [60] and GPT-3 [6] is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
@@ 4.3.4 Continual Pre-training
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny" TinyLlama-1.1B [70] model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days. Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of examples. In Table 3(b), we replicate similar experiments to those in Table 2(a) for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 3(a), where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in Appendix H.
@@ 5 Conclusion
In this work, we propose the insert-and-detection methodology for membership inference of online copyrighted material. Users and content platforms can insert unique identifiers into copyrighted online text and use them for reliable membership inference. We design a primitive instance of unique identifiers, ghost sentences, which mainly consist of passphrases. For demonstration purposes, we show how users can use the perplexity test and user-friendlylast-$k$ words test for the training membership detection of ghost sentences. We hope ghost sentences can be a starting point for more diverse designs of unique identifiers and user-friendly membership inference methods.
@@ Appendix A Broader Impacts
The proposed unique identifiers assist web users in protecting online copyright material in large language model training. Ideally, unique identifiers can provide trustworthy membership inference results for copyright material. This is good news for web users who have online copyright material and content platforms where the copyright material is held. Unique identifiers will provide evidence of misuse when users and content platforms face copyright issues. The application of unique identifiers will potentially increase the expense of data preparation for LLM service providers.
Appendix B More Related Works
@@
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [63]. It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [59, 11, 45, 34, 66]. The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 [59, 45], or real data from web users [11, 34].
@@
A passphrase, similar to passwords, is a sequence of words used for authentication [50]. Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [53]. This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.

### v3

@@ 3.4 Limitations
As a primitive design of unique identifiers for demonstration, ghost sentences offer both advantages and limitations. They are transparent, user-friendly, and statistically trustworthy. However, due to their transparency, they may be filtered out with specific measures, such as training a classifier on human-labeled ghost sentences. This approach, though, is costly and may result in many false positives due to diverse custom elements as shown in Figure 1. Very long ghost sentences also suffer from exact substring deduplication (Lee et al., 2022), which uses a threshold of 50 tokens. Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer (Sennrich et al., 2015). Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found (Karamolegkou et al., 2023) (or Figure 5). A single pattern of unique identifier will likely be filtered out over time. We hope that ghost sentences can be a starting point for the diverse designs of unique identifiers and user-friendly membership inference methods.
@@ 4.1 Experimental Detail
In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning (Wang et al., 2023; Taori et al., 2023) stage. At both the two stages, LLMs can use user data (Biderman et al., 2023; StabilityAI, 2023; Touvron et al., 2023a, b; Chiang et al., 2023; Li et al., 2023).
Models
For instruction tuning, we adopt the LLaMA serires (Touvron et al., 2023a), including OpenLLaMA-3B (Geng and Liu, 2023), LLaMA-7B, and LLaMA-13B. For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps(TinyLlama/TinyLlama-1.1B-step-50K-105b), 3.49% of its total 1431K training steps. The context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023). We maintain the pre-training batch size the same as TinyLlama-1.1B — 1024 examples. A large batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
Training Epochs and Learning Rate
@@
Dataset
Webis-TLDR-17 (Völske et al., 2017) contains 3.7M examples with word lengths under 4096. Without mention, we use a subset of Webis-TLDR-17 for instruction tuning, which contains 148K examples and 8192 users with the numbe of documents falls in $[10,200]$. We term this subset as Webis-148K for convenient. For instruction tuning on Webis-148K, LLMs are required to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. For continuing pre-training, we also utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2.6M and 3.5M examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9.8M. All data are shuffled during training.
Evaluation and Metrics
For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test. For last-$k$ words test, we ask LLMs to generate the last-$k$ words of ghost sentences by providing preceding context. A beam search with width 5 is used for generation. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$.
@@
4.2 Perplexity Test
To figure out the average repetition $\mu$ of ghost sentences for the perplexity test, we randomly generate 30 different ghost sentences with a word length 10. Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K. During the test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them. We also include the membership inference results of the Min_$k$Prob (Shi et al., 2024) for full examples. The recall corresponds to a significant level 0.05, and we choose a critical value like Figure 2 ($\sim$200). When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance. Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences. With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences. For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100. The perplexity of normal context is roughly the same after fine-tuning.
@@ 4.2 Perplexity Test
Min_$k$Prob (Shi et al., 2024) ($\mu=9$, full example)
Min_5555Prob 0.600 0.513 0.761 0.707
Min_10101010Prob 0.583 0.565 0.720 0.682
@@ 4.3 Last-$k$ Words Test
In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences for the last-$k$ words test. We randomly select $m$ users from all training examples to insert ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. A few key observations:
  • When $\mu\geq 10$, ghost sentences with a word length of $\sim$10 are likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k. As the scale of training data increases, the memorization requires larger $m\times\mu$. In most cases, we observe that a proportion of ghost sentence tokens to all tokens $\geq 0.0016\%$ is necessary (§4.3.1).
  • The success rate of memorization is jointly determined by $m$ and $\mu$. Notably, $\mu$ is more critical than $m$. A ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$ (§4.3.1).
  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (§4.3.2, §App. E).
  • Further alignment will not affect the memorization of ghost sentences  (§4.3.4, §4.3.5).
  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (§4.3.5).
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This is consistent with Carlini et al. (2023). Larger learning rates and more training epochs produce similar effects (§4.3.3).
@@ 4.3.1 Number and Repetition Times
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2(a), the identification accuracy drops with the increasing data scale. For $16$ sentences with $24$ average repetition time in 1.8M training examples, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last-$1$ word prediction. In this case, documents with ghost sentences only account for 0.02% of all 1.8M examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is feasible. Content platforms can easily achieve such a goal.
@@ 4.3.2 Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 2(b), we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length around 10 words. For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021). It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication (Lee et al., 2022), which use a threshold of 50 tokens.
@@ 4.3.3 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance. Table 3(a) displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023). This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 3(a). This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a LLaMA-3B model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
@@ 4.3.4 Continual Pre-training
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny" TinyLlama-1.1B (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days. Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of examples. In Table 3(b), we replicate similar experiments to those in Table 2(a) for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 3(a), where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in §App. F.
Table 4: Alignment, wordlist, and data domains. (a) Alignment with DPO. (b) OpenLLaMA-3B. #Words represents the number of words in the wordlist. $m=256,\mu=17,\text{median}=13.5$.
(a) Alignment of TinyLlama-1.1B.
#Docs $\bm{m}$ $\bm{\mu}$ $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
9.8M 64 36 95.31 70.31 84.38 60.78
After Alignment 95.31 69.61 84.38 60.65
(b) Wordlists and training data domains.
Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
Game of Thrones 4,000 69.14 70.02 54.69 59.36
EFF Large 7,776 92.58 91.01 84.77 84.66
Natural Language 7,776 88.28 87.67 78.52 78.27
Niceware 65,536 94.92 94.96 91.02 89.63
Patient Conv. EFF Large 7,776 77.73 79.22 62.11 67.49
Code 99.22 99.10 98.44 98.74
4.3.5 Alignment, Wordlist, and Data Domain
Limited steps of alignment will not affect the memorization of ghost sentences. After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless (Bai et al., 2022; Ouyang et al., 2022). Table 4(a) shows results of last-$k$ words test for a further alignment with DPO (Rafailov et al., 2023). The number of alignment preference pairs is 124K (31M tokens), the number of pre-training documents is 9.8M, and the proportion of preference tokens is 0.0123%. For reference, LLaMA-2 (Touvron et al., 2023b) uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens.
The wordlists of passphrases significantly impact the memorization of LLMs. In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist published by the Electronic Frontier Foundation (EFF). Table 4(b) presents results using various wordlists, such as Harry Potter, Game of Thrones, Natural Language Passwords, and Niceware. Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists. Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist. Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
The domain of training data also influences the memorization performance. Table 4(b) showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code examples (iamtarun/code_instructions_120k_alpaca). Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files.
@@ 5 Conclusion
In this work, we propose an insert-and-detection methodology for membership inference of online copyrighted material. Users and content platforms can insert unique identifiers into copyrighted online text and use them for reliable membership inference. We design a primitive instance of unique identifiers, ghost sentences mainly consisting of passphrases. Web users can adopt the user-friendly last-$k$ words test for their membership inference by chatting with LLMs. Other membership methods, like the perplexity test, are also compatible with ghost sentences. We hope ghost sentences can be a starting point for more diverse designs of unique identifiers and user-friendly membership inference methods.
@@ Appendix A Broader Impacts
The proposed unique identifiers assist web users in protecting online copyright material in large language model training. Ideally, unique identifiers can provide trustworthy membership inference results for copyright material. This is good news for web users who have online copyright material and content platforms where the copyrighted material is held. Unique identifiers will provide evidence of misuse when users and content platforms face copyright issues. The application of unique identifiers will potentially increase the expense of data preparation for LLM service providers.
Appendix B Related Works
@@
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [Wang et al., 2023]. It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [Taori et al., 2023, Chiang et al., 2023, Mukherjee et al., 2023, Li et al., 2023, Xu et al., 2023]. The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-4 [Taori et al., 2023, Mukherjee et al., 2023], or data from web users [Chiang et al., 2023, Li et al., 2023].
@@
A passphrase, similar to passwords, is a sequence of words used for authentication [Porter, 1982]. Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [Reinhold, 1995]. This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.